OpenStudy (anonymous):

if someone can help me out with figuring out how to solve this using substitution that would be awesome my teacher doesn't know how to teach or help Solve the systems using substitution. 6x + 2y = 0 4x + y = -1

OpenStudy (anonymous):

y = -4x-1 so 6x + -4x-1= 0

OpenStudy (anonymous):

basically solve for one letter, then plug it in the other question so that one equation only has one type of letter. then solve for that letter, then plug it back in the other equation to find the other letter.

OpenStudy (anonymous):

hah i dont uknow dis stuff but nice pic. by the way

OpenStudy (anonymous):

thanks

OpenStudy (texaschic101):

4x + y = -1 -- we want to try to get a variable by itself and this equation looks the easiest....subtract 4x from both sides y = -4x - 1 now we sub -4x - 1 in for y in the other equation 6x + 2y = 0 6x + 2(-4x - 1) = 0 --- distribute through the parenthesis 6x - 8x - 2 = 0 -- add 2 to both sides 6x - 8x = 2 + 0 -- combine like terms -2x = 2 -- divide both sides by -2 x = -1 now sub -1 in for x in either of your original equations to find y 4x + y = -1 4(-1) + y = -1 -4 + y = -1 y = -1 + 4 y = 3 check your answers.. 4x + y = -1 4(-1) + 3 = -1 -4 + 3 = -1 -1 = -1 (correct) so x = -1 and y = 3

OpenStudy (anonymous):

lol not you but yeah i guess you too ^_^ but i meant @austindrury

OpenStudy (anonymous):

my pic is the best

OpenStudy (anonymous):

well thats affensive

OpenStudy (anonymous):

yeah ikr :D

OpenStudy (anonymous):

well I took calculus not dumb

OpenStudy (anonymous):

:3

OpenStudy (anonymous):

@got1that1gas1money you've been BURNED!!! hahahahaha XD

OpenStudy (anonymous):

you want some ice with that burn xD go @RyleeC!!!!!

OpenStudy (anonymous):

Told someone to stop yelling in my face before mine gets infectedxD

OpenStudy (anonymous):

Yeah im not so good with combacks:P im horrible

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

hahahaha wonderful i love it ^_^ thank you by the way @RainbowBubbles01 i like your picture as well its gorgeous and trippy :D i enjoy trippy

OpenStudy (anonymous):

your welcome and thaqnk you so much no one has commented on my picture so far:D

OpenStudy (anonymous):

thank*

OpenStudy (anonymous):

mine either! this is the first time i logged on here and you made it dope!!!! thank you all for the feedback @texaschic101 @RyleeC @RainbowBubbles01 and @got1that1gas1money go tend your burns homie

OpenStudy (anonymous):

lmao ^_^

OpenStudy (anonymous):

for sure

OpenStudy (anonymous):

;)

OpenStudy (anonymous):

everyone tell two of your favorite yo momma joke xD sees whos is better

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

ok iguess wrong time

OpenStudy (anonymous):

yo mama so stupid when she threw a rock at the ground she missed

OpenStudy (anonymous):

yo mama so fat when she sat on the iphone she made the i pad @RainbowBubbles01

OpenStudy (anonymous):

lol your second 1?@austindrury

OpenStudy (anonymous):

my first one is Yo momma so ugly when one direction saw her they went the other direction

OpenStudy (anonymous):

and of course my second one is yo momma so old when she tries the breast feed her milk is powdered milk

OpenStudy (anonymous):

hahahahaha XD i got another one

OpenStudy (anonymous):

ok lets see it

OpenStudy (anonymous):

do you play combat arms?

OpenStudy (anonymous):

You mama so fat! her belly button gets home 15 minutes before she does

OpenStudy (anonymous):

combat arms is that a facebook game??

OpenStudy (anonymous):

nother one!!!

OpenStudy (anonymous):

your mamas so fat that the national weather service names each one of her fartsXD

OpenStudy (anonymous):

lol no you have to download it i dont like facebook games

OpenStudy (anonymous):

ill check it out i dont have a iphone or anything

OpenStudy (anonymous):

i didnt say anything about a iphone or anything xD

OpenStudy (anonymous):

yo mamas so fat she threw on a sheet for halloween and went as antarctica. haha - Austins girlfriend

OpenStudy (anonymous):

haha where do i download it??

OpenStudy (anonymous):

online look up combat arms fps game and download it tell me when you do an dmy username is ki11rbunny01

OpenStudy (anonymous):

have you looked it up?

OpenStudy (anonymous):

ki11rbunny01

OpenStudy (anonymous):

and i can help you level up easily and you should go through the training or else you wont understand how to play

OpenStudy (anonymous):

hellur?

OpenStudy (anonymous):

hell yeah im at school right now so ill check it out later and add you

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

did you write my username down

OpenStudy (anonymous):

and what will yours be so i know its you

OpenStudy (anonymous):

yeah i did mine will probably be reeferkeifer

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

mine has no caps. oor spaces

OpenStudy (anonymous):

ki11rbunny01

OpenStudy (anonymous):

and yes im a girl

OpenStudy (anonymous):

and iplay a gun game sowhat

OpenStudy (anonymous):

no caps or spaces got it ^_^ hahaha thats awsome

OpenStudy (anonymous):

ikr adn i love yours ^_^ :3

OpenStudy (anonymous):

/) /) ( ^_^) C (") (")

OpenStudy (anonymous):

bunny O_O

OpenStudy (anonymous):

thank you :D that bunny is amazing

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

tell me when its dowloaded